A falling body has an initial velocity of 200cm/sec.How far will it drop before its velocity becomes 10,000cm/sec


1
Expert's answer
2021-01-31T03:40:51-0500

The final velocity

"v_f=v_i+gt"

Hence

"t=\frac{v_f-v_i}{g}=\frac{100\:\rm m/s-2\:\rm m/s}{9.8\:\rm m/s^2}=10\:\rm s"